ata high school basketball game, 400 tickets are sold. adult tickets cost 5 dollars and student tickets cost 2.50. if the totla amount collected was 1375 how many student tickets were sold?
Found 2 solutions by rfer, unlockmath:
Answer by rfer(16322)

a+s=400 a=400-s
5a+2.5s=1375
-----------
5(400-s)+2.5s=1375
2000-5s+2.5s=1375
-2.5s=-625
s=250 students
a=150 adults

Answer by unlockmath(1688)

Hello,
We can set up two equations. Let x be the adult tickets and y for the student tickets.
5x+2.5y=1375
x+y=400
Now multiply the second equation by -5 to get:
-5x-5y=-2000 Now add this to the other equation:
5x+2.5=1375 This we give us:
-2.5y=-625 Divide both sides by -2.5 to get:
y=250
There we go 250 student tickets were sold.
You can plug in the answer and see if it works.
Make sense?
RJ
You might check out a book I wrote at:
www.math-unlock.com